# Gravity and fluid dynamic correspondence on a null hypersurface: inconsistencies and advancement

### v1

@@ Acknowledgement
The research of SD is supported by the Polish National Science Centre, as part of the OPUS 22 project number 2021/43/B/ST2/02950. The research of one of the authors (BRM) is supported by the Science and Engineering Research Board (SERB), Department of Science &\&& Technology (DST), Government of India, under the scheme Core Research Grant (File no. CRG/2020/000616).
@@
Appendix A Derivation of Eq. 37
The conjugate momentum $\mathcal{P}_{a}$, given in 36, upon use of $\omega_{a}=\Omega_{a}-\kappa k_{a}$ can be expressed as
@@
On the other hand using (36) and $\theta=-(1/2)q_{ij}\pounds_{l}q^{ij}$ we have
@@
Finally, combining A.4 and A.5 we find 37.
@@ Appendix B Conservation of null BY tensor
The expression of null BY tensor is given by 39. The differentiation operator $\mathcal{D}_{a}$ operates on contravariant and covariant vectors as $\mathcal{D}_{a}A^{b}=\Pi^{i}_{~{}a}\Pi^{b}_{~{}j}\nabla_{i}(\Pi^{j}_{~{}k}A^{k})$ and $\mathcal{D}_{a}B_{b}=\Pi^{i}_{~{}a}\Pi^{j}_{~{}b}\nabla_{i}(\Pi^{k}_{~{}j}B_{k})$, which we have mentioned earlier. Therefore, we obtain
$\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=\Pi^{i}_{~{}a}\Pi^{a}_{~{}j}\Pi^{k}_{~{}c}\nabla_{i}T^{j}_{~{}~{}k}|_{(null)}=\Pi^{i}_{~{}j}\Pi^{k}_{~{}c}\nabla_{i}T^{j}_{~{}~{}k}|_{(null)}~{},$ (B.1)
@@
$\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=\Pi^{b}_{~{}c}\Big{(}\nabla_{a}T^{a}_{~{}~{}b}|_{(null)}-k^{i}\nabla_{i}l_{a}T^{a}_{~{}~{}b}|_{(null)}\Big{)}~{},$ (B.2)
using the explicit expression of the null BY tensor, as provided in 39, one obtains
$\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=\frac{1}{8\pi}\Pi^{b}_{~{}c}\Big{[}\nabla_{a}\theta^{a}_{b}+\Big{(}\theta+\kappa\Big{)}\omega_{b}+l^{a}\nabla_{a}\omega_{b}-\Big{(}\theta+\kappa\Big{)}k^{a}\nabla_{a}l_{b}-\Pi^{a}_{~{}b}\nabla_{a}\Big{(}\theta+\kappa\Big{)}\Big{]}$
@@
$\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=\frac{1}{8\pi}\Pi^{b}_{~{}c}\Big{[}\nabla_{a}\theta^{a}_{b}+\Big{(}\theta+\kappa\Big{)}\omega_{b}+l^{a}\nabla_{a}\omega_{b}-\Pi^{a}_{~{}b}\nabla_{a}\Big{(}\theta+\kappa\Big{)}-\theta^{a}_{b}(k^{i}\nabla_{i}l_{a})\Big{]}$ (B.4)
@@
$R_{ab}l^{b}=\nabla_{m}\theta^{m}_{a}+l^{m}\nabla_{m}\omega_{a}+\Big{(}\theta+\kappa\Big{)}\omega_{a}-\nabla_{a}\Big{(}\theta+\kappa\Big{)}-\theta_{am}k^{n}\nabla_{n}l^{m}$
$-\Big{(}\omega_{m}k^{n}\nabla_{n}l^{m}+\nabla_{m}k^{n}\nabla_{n}l^{m}+k^{n}\nabla_{m}\nabla_{n}l^{m}\Big{)}l^{a}~{}.$ (B.5)
@@
$R_{ab}l^{b}\Pi^{a}_{~{}c}=\Pi^{a}_{~{}c}\Big{[}\nabla_{m}\theta^{m}_{a}+l^{m}\nabla_{m}\omega_{a}+\Big{(}\theta+\kappa\Big{)}\omega_{a}-\nabla_{a}\Big{(}\theta+\kappa\Big{)}-\theta_{am}k^{n}\nabla_{n}l^{m}\Big{]}~{}.$ (B.6)
Combining B.4 and B.6, we obtain 41.
Appendix C Energy conservation relation in null hydrodynamics

### v2

@@ Acknowledgement
The research of KB is supported by the New Faculty Seed Grant (NFSG) of BITS Pilani Dubai Campus. The research of SD is supported by the Polish National Science Centre, as part of the OPUS 22 project number 2021/43/B/ST2/02950. The research of BRM is supported by the Science and Engineering Research Board (SERB), Department of Science &\& Technology (DST), Government of India, under the scheme Core Research Grant (File no. CRG/2020/000616).
@@
Appendix A Derivation of Eq. Eq. (
The conjugate momentum $\mathcal{P}_{a}$, given in Eq. (, upon use of $\omega_{a}=\Omega_{a}-\kappa k_{a}$ can be expressed as
@@
On the other hand using (Eq. () and $\theta=-(1/2)q_{ij}\pounds_{l}q^{ij}$ we have
@@
Finally, combining Eq. ( and Eq. ( we find Eq. (.
@@ Appendix B Conservation of null BY tensor
The expression of null BY tensor is given by Eq. (. The differentiation operator $\mathcal{D}_{a}$ operates on contravariant and covariant vectors as $\mathcal{D}_{a}A^{b}=\Pi^{i}_{~a}\Pi^{b}_{~j}\nabla_{i}(\Pi^{j}_{~k}A^{k})$ and $\mathcal{D}_{a}B_{b}=\Pi^{i}_{~a}\Pi^{j}_{~b}\nabla_{i}(\Pi^{k}_{~j}B_{k})$, which we have mentioned earlier. Therefore, we obtain
$\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=\Pi^{i}_{~a}\Pi^{a}_{~j}\Pi^{k}_{~c}\nabla_{i}T^{j}_{~~k}|_{\rm(null)}=\Pi^{i}_{~j}\Pi^{k}_{~c}\nabla_{i}T^{j}_{~~k}|_{\rm(null)}~,$ (B.1)
@@
$\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=\Pi^{b}_{~c}\Big(\nabla_{a}T^{a}_{~~b}|_{\rm(null)}-k^{i}\nabla_{i}l_{a}T^{a}_{~~b}|_{\rm(null)}\Big)~,$ (B.2)
using the explicit expression of the null BY tensor, as provided in Eq. (, one obtains
$\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=\frac{1}{8\pi}\Pi^{b}_{~c}\Big[\nabla_{a}\theta^{a}_{b}+\Big(\theta+\kappa\Big)\omega_{b}+l^{a}\nabla_{a}\omega_{b}-\Big(\theta+\kappa\Big)k^{a}\nabla_{a}l_{b}-\Pi^{a}_{~b}\nabla_{a}\Big(\theta+\kappa\Big)\Big]$
@@
$\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=\frac{1}{8\pi}\Pi^{b}_{~c}\Big[\nabla_{a}\theta^{a}_{b}+\Big(\theta+\kappa\Big)\omega_{b}+l^{a}\nabla_{a}\omega_{b}-\Pi^{a}_{~b}\nabla_{a}\Big(\theta+\kappa\Big)-\theta^{a}_{b}(k^{i}\nabla_{i}l_{a})\Big]$ (B.4)
@@
$R_{ab}l^{b}=\nabla_{m}\theta^{m}_{a}+l^{m}\nabla_{m}\omega_{a}+\Big(\theta+\kappa\Big)\omega_{a}-\nabla_{a}\Big(\theta+\kappa\Big)-\theta_{am}k^{\rm N}\nabla_{n}l^{m}$
$-\Big(\omega_{m}k^{\rm N}\nabla_{n}l^{m}+\nabla_{m}k^{\rm N}\nabla_{n}l^{m}+k^{\rm N}\nabla_{m}\nabla_{n}l^{m}\Big)l^{a}~.$ (B.5)
@@
$R_{ab}l^{b}\Pi^{a}_{~c}=\Pi^{a}_{~c}\Big[\nabla_{m}\theta^{m}_{a}+l^{m}\nabla_{m}\omega_{a}+\Big(\theta+\kappa\Big)\omega_{a}-\nabla_{a}\Big(\theta+\kappa\Big)-\theta_{am}k^{\rm N}\nabla_{n}l^{m}\Big]~.$ (B.6)
Combining Eq. ( and Eq. (, we obtain Eq. (.
Appendix C Some specific examples
C.1 Schwarzschild spacetime
To illustrate the general formalism, we consider the Schwarzschild spacetime and explicitly compute the null Brown–York tensor and the associated fluid variables. In ingoing Eddington–Finkelstein coordinates $(t,r,\psi,\phi)$, the Schwarzschild metric is given by
$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{4M}{r}\,dt\,dr+\left(1+\frac{2M}{r}\right)dr^{2}+r^{2}\left(d\psi^{2}+\sin^{2}\psi\,d\phi^{2}\right).$ (C.1)
A convenient choice of outgoing null generator and its dual one-form is
[…]
The auxiliary ingoing null vector field is chosen as
[…]
With this choice, the normalization conditions
$l^{a}l_{a}=0=k^{a}k_{a},\qquad l^{a}k_{a}=-1$ (C.4)
are easily verified.
The mixed projection tensor $\Pi^{a}{}_{b}=\delta^{a}{}_{b}+k^{a}l_{b}$ takes the form
$\Pi^{a}{}_{b}=\begin{pmatrix}\frac{M}{r}+\frac{1}{2}&\frac{M}{r}+\frac{1}{2}&0&0\\ \frac{1}{2}-\frac{M}{r}&\frac{1}{2}-\frac{M}{r}&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$. (C.5)
The induced metric on the two-dimensional spatial cross-section $\mathcal{B}_{(\text{null})}$ is
$q_{ab}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&r^{2}&0\\ 0&0&0&r^{2}\sin^{2}\psi\end{pmatrix}$, (C.6)
which correctly reproduces the metric on a round two-sphere of radius $r$.
The nonvanishing components of the second fundamental form $\theta_{ab}$ are found to be
$\theta_{ab}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&\dfrac{r(r-2M)}{2M+r}&0\\ 0&0&0&-\dfrac{r\sin^{2}\psi(2M-r)}{2M+r}\end{pmatrix}$. (C.7)
Taking the trace with respect to $q^{ab}$ yields the expansion scalar
$\theta=\frac{2(r-2M)}{r(2M+r)}.$ (C.8)
As expected for a spherically symmetric spacetime, the shear tensor vanishes identically,
$\sigma_{ab}=0.$ (C.9)
The non-affinity parameter (surface gravity) associated with the null generators is
$\kappa=\frac{4M}{(2M+r)^{2}}.$ (C.10)
Using these quantities, the null Brown–York–like energy–momentum tensor defined in Eq. ( evaluates to
$T^{a}{}_{b}\big|_{\rm(null)}=\begin{pmatrix}-\dfrac{r-2M}{8\pi r^{2}}&-\dfrac{r-2M}{8\pi r^{2}}&0&0\\ -\dfrac{(r-2M)^{2}}{8\pi r^{2}(2M+r)}&-\dfrac{(r-2M)^{2}}{8\pi r^{2}(2M+r)}&0&0\\ 0&0&-\dfrac{-4M^{2}+4Mr+r^{2}}{8\pi r(2M+r)^{2}}&0\\ 0&0&0&-\dfrac{-4M^{2}+4Mr+r^{2}}{8\pi r(2M+r)^{2}}\end{pmatrix}$. (C.11)
The effective fluid variables on the null surface can now be extracted in a straightforward manner. The quasilocal energy density is
$\epsilon_{\rm null}=-T^{a}{}_{b}\big|_{\rm(null)}\,k_{a}l^{b}=\frac{2M-r}{8\pi Mr+4\pi r^{2}},$ (C.12)
which is negative outside the horizon $(r>2M)$, consistent with the attractive nature of gravity. The momentum density vanishes identically,
$p^{c}=-T^{a}{}_{b}\big|_{\rm(null)}\,k_{a}q^{bc}=0,$ (C.13)
reflecting the absence of rotation or angular momentum flux in the Schwarzschild geometry.
Finally, the spatial stress tensor on $\mathcal{B}_{(\text{null})}$ is given by
$s^{ab}=q^{a}{}_{c}q^{bd}T^{c}{}_{d}\big|_{\rm(null)}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&-\dfrac{-4M^{2}+4Mr+r^{2}}{8\pi r^{3}(2M+r)^{2}}&0\\ 0&0&0&-\dfrac{\csc^{2}\psi(-4M^{2}+4Mr+r^{2})}{8\pi r^{3}(2M+r)^{2}}\end{pmatrix}$. (C.14)
This completes the explicit realization of the null fluid variables for the Schwarzschild spacetime and provides a concrete consistency check of our general formalism.
C.2 Kerr spacetime
We now extend the null Brown–York construction to the Kerr spacetime, describing a stationary, axisymmetric rotating black hole of mass $M$ and angular momentum $J$. In Boyer–Lindquist coordinates $(t,r,\psi,\phi)$, the Kerr metric takes the form
$ds^{2}=$ $-\left(1-\frac{2Mr}{\Sigma}\right)dt^{2}+\frac{\Sigma}{\Delta}\,dr^{2}+\Sigma\,d\psi^{2}-\frac{4aMr\sin^{2}\psi}{\Sigma}\,dt\,d\phi$
$+\sin^{2}\psi\left(r^{2}+a^{2}+\frac{2a^{2}Mr\sin^{2}\psi}{\Sigma}\right)d\phi^{2},$ (C.15)
where
$\Sigma$ $=r^{2}+a^{2}\cos^{2}\psi,$ (C.16)
$\Delta$ $=r^{2}-2Mr+a^{2},$ (C.17)
and $a\equiv J/M$ denotes the specific angular momentum.
To construct the null boundary geometry, we choose a pair of future-directed null vectors $l^{a}$ and $k^{a}$, satisfying $l^{a}l_{a}=k^{a}k_{a}=0$ and $l^{a}k_{a}=-1$. A convenient choice adapted to the Kerr geometry is
$l^{a}=\left(\frac{r^{2}+a^{2}}{\Delta},\;1,\;0,\;\frac{a}{\Delta}\right),\qquad l_{a}=\left(-1,\;\frac{\Sigma}{\Delta},\;0,\;a\sin^{2}\psi\right),$ (C.18)
together with
$k^{a}=\left(\frac{r^{2}+a^{2}}{2\Sigma},-\frac{\Delta}{2\Sigma},0,\frac{a}{2\Sigma}\right),\qquad k_{a}=\left(-\frac{\Delta}{2\Sigma},\;-\frac{1}{2},\;0,\;\frac{a\,\Delta\,\sin^{2}\psi}{2\Sigma}\right).$ (C.19)
This choice ensures an affine parametrization of the null generators, as confirmed by the vanishing non-affinity parameter $\kappa=0$.
The mixed projector orthogonal to the null pair is defined as $\Pi^{a}{}_{b}=\delta^{a}{}_{b}+k^{a}l_{b}$, which evaluates to
$\Pi^{a}{}_{b}=\begin{pmatrix}1-\dfrac{r^{2}+a^{2}}{2\Sigma}&\dfrac{r^{2}+a^{2}}{2\Delta}&0&\dfrac{a(r^{2}+a^{2})\sin^{2}\psi}{2\Sigma}\\[6.0pt] \dfrac{\Delta}{2\Sigma}&\dfrac{1}{2}&0&-\dfrac{a\Delta\sin^{2}\psi}{2\Sigma}\\[6.0pt] 0&0&1&0\\[6.0pt] -\dfrac{a}{2\Sigma}&\dfrac{a}{2\Delta}&0&1+\dfrac{a^{2}\sin^{2}\psi}{2\Sigma}\end{pmatrix}$. (C.20)
The induced metric on the spatial cross-sections of the null hypersurface, $q_{ab}=g_{ab}+k_{a}l_{b}+l_{a}k_{b}$, is given by
$q_{ab}=\begin{pmatrix}\dfrac{a^{2}\sin^{2}\psi}{\Sigma}&0&0&-\dfrac{a(r^{2}+a^{2})\sin^{2}\psi}{\Sigma}\\[6.0pt] 0&0&0&0\\[6.0pt] 0&0&\Sigma&0\\[6.0pt] -\dfrac{a(r^{2}+a^{2})\sin^{2}\psi}{\Sigma}&0&0&\dfrac{(r^{2}+a^{2})^{2}\sin^{2}\psi}{\Sigma}\end{pmatrix}$. (C.21)
The null extrinsic curvature $\theta_{ab}$, encoding the expansion and shear of the null congruence, is found to be
$\theta_{ab}=\begin{pmatrix}\dfrac{a^{2}r\sin^{2}\psi}{\Sigma^{2}}&0&\dfrac{a^{2}\sin 2\psi}{2\Sigma}&-\dfrac{ar(r^{2}+a^{2})\sin^{2}\psi}{\Sigma^{2}}\\[6.0pt] 0&0&0&0\\[6.0pt] -\dfrac{a^{2}\sin 2\psi}{2\Sigma}&0&r&\dfrac{a(r^{2}+a^{2})\sin 2\psi}{2\Sigma}\\[6.0pt] -\dfrac{ar(r^{2}+a^{2})\sin^{2}\psi}{\Sigma^{2}}&0&-\dfrac{a(r^{2}+a^{2})\sin 2\psi}{2\Sigma}&\dfrac{r(r^{2}+a^{2})^{2}\sin^{2}\psi}{\Sigma^{2}}\end{pmatrix}$. (C.22)
Its trace yields the null expansion
$\theta=\frac{2r}{\Sigma},$ (C.23)
We emphasize that the expansion scalar depends on the choice and normalization of the null generator. The result $\theta=2r/\Sigma$ reduces to $\theta=2/r$ in the limit $a\to 0$, corresponding to the principal null congruence of Schwarzschild spacetime. This differs from the Eddington–Finkelstein result obtained earlier, which is adapted to horizon-regular null generators.
The traceless part of $\theta_{ab}$ defines the shear tensor,
$\sigma_{ab}=\begin{pmatrix}0&0&\dfrac{a^{2}\sin 2\psi}{2\Sigma}&0\\[6.0pt] 0&0&0&0\\[6.0pt] -\dfrac{a^{2}\sin 2\psi}{2\Sigma}&0&0&\dfrac{a(r^{2}+a^{2})\sin 2\psi}{2\Sigma}\\[6.0pt] 0&0&-\dfrac{a(r^{2}+a^{2})\sin 2\psi}{2\Sigma}&0\end{pmatrix}$, (C.24)
explicitly capturing anisotropic distortions induced by rotation.
The null Brown–York stress tensor on the null boundary is obtained as
$T^{a}{}_{b}\big|_{\rm(null)}=\frac{1}{8\pi}\begin{pmatrix}-\dfrac{r}{\Sigma}&-\dfrac{r(r^{2}+a^{2})}{\Delta\,\Sigma}&-\dfrac{a^{2}\sin 2\psi}{\Sigma}&-\dfrac{ar(r^{2}+a^{2})\sin^{2}\psi}{\Delta\,\Sigma}\\[8.0pt] -\dfrac{r\Delta}{\Sigma^{2}}&-\dfrac{r}{\Sigma}&-\dfrac{a^{2}\sin 2\psi}{2\Sigma}&-\dfrac{ar\sin^{2}\psi}{\Sigma}\\[8.0pt] -\dfrac{a^{2}\sin 2\psi}{\Sigma^{2}}&0&-\dfrac{r}{\Sigma}&\dfrac{a(r^{2}+a^{2})\sin 2\psi}{\Sigma^{2}}\\[8.0pt] \dfrac{ar\sin^{2}\psi}{\Sigma^{2}}&-\dfrac{ar}{\Delta\,\Sigma}&\dfrac{a\cot\psi}{2\Sigma}&-\dfrac{r(r^{2}+a^{2})\sin^{2}\psi}{\Delta\,\Sigma}\end{pmatrix}$. (C.25)
From this tensor, the null fluid variables are readily identified. The energy density measured by the null observer is
$\epsilon_{\rm null}=-\frac{r}{4\pi\,\Sigma},$ (C.26)
which reduces to $\epsilon_{\rm null}=-1/(4\pi r)$ in the limit $a\to 0$, corresponding to the principal null congruence of Schwarzschild spacetime. The momentum density on the spatial cross-sections is
$p^{c}=\left(-\frac{a^{2}r\sin^{2}\psi}{8\pi\Sigma^{2}},\;0,\;-\frac{a^{2}\sin 2\psi}{16\pi\Sigma^{2}},\;-\frac{ar}{8\pi\Sigma^{2}}\right),$ (C.27)
vanishing identically in the non-rotating limit.
Finally, the spatial stress tensor takes the form
$s^{ab}=\begin{pmatrix}-\dfrac{a^{2}r\sin^{2}\psi}{8\pi\Sigma^{2}}&0&-\dfrac{a^{2}\sin 2\psi}{16\pi\Sigma^{2}}&-\dfrac{ar}{8\pi\Sigma^{2}}\\[8.0pt] 0&0&0&0\\[8.0pt] \dfrac{a^{2}\sin 2\psi}{16\pi\Sigma^{2}}&0&-\dfrac{r}{8\pi\Sigma^{2}}&\dfrac{a\cot\psi}{8\pi\Sigma^{2}}\\[8.0pt] -\dfrac{ar}{8\pi\Sigma^{2}}&0&-\dfrac{a\cot\psi}{8\pi\Sigma^{2}}&-\dfrac{r\csc^{2}\psi}{8\pi\Sigma^{2}}\end{pmatrix}$, (C.28)
demonstrating explicitly how rotation induces anisotropic stresses on the null surface.
In summary, while the Schwarzschild spacetime leads to a remarkably simple null fluid description characterized by vanishing shear, zero momentum density, and isotropic spatial stress, the Kerr spacetime exhibits a much richer structure induced by rotation. The presence of angular momentum introduces non-vanishing shear, anisotropic stresses, and a finite momentum density on the null hypersurface, reflecting the coupling between geometry and frame dragging. Nevertheless, the expansion scalar retains a universal form, reducing smoothly to the Schwarzschild result in the non-rotating limit. This comparison clearly illustrates how rotation imprints itself on the null Brown–York tensor, providing a clean geometric manifestation of rotational degrees of freedom within the null fluid framework.
C.3 FLRW metric
Having analyzed stationary black hole spacetimes, it is instructive to briefly apply our formalism to a cosmological setting. We consider the spatially flat Friedmann–Robertson–Walker (FRW) spacetime, which describes a homogeneous and isotropic expanding universe. In spherical polar coordinates, the metric is given by
$ds^{2}=-dt^{2}+a^{2}(t)\left[dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right)\right]~,$ (C.29)
where $a(t)$ is the cosmological scale factor. Radial null generators adapted to the cosmological symmetry can be chosen as
$l^{a}=\left(1,\;\frac{1}{a(t)},\;0,\;0\right),\qquad l_{a}=\left(-1,\;a(t),\;0,\;0\right),$ (C.30)
together with the auxiliary null vector
$k^{a}=\left(\frac{1}{2},\,-\frac{1}{2a(t)},\,0,\,0\right),\qquad k_{a}=\left(-\frac{1}{2},\,-\frac{a(t)}{2},\,0,\,0\right),$ (C.31)
which satisfy the normalization conditions $l^{a}l_{a}=0=k^{a}k_{a}$ and $l^{a}k_{a}=-1$.
The mixed projection tensor orthogonal to the null generators is given by
$\Pi^{a}{}_{b}=\left(\begin{array}[]{cccc}\frac{1}{2}&\frac{a(t)}{2}&0&0\\ \frac{1}{2a(t)}&\frac{1}{2}&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right),$ (C.36)
and the induced metric on the two-dimensional spatial cross-sections takes the form
$q_{ab}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&r^{2}a^{2}(t)&0\\ 0&0&0&r^{2}a^{2}(t)\sin^{2}\theta\end{array}\right).$ (C.41)
As expected, the induced geometry corresponds to a round two-sphere of areal radius $R=a(t)r$. The non-vanishing components of the second fundamental form are
$\theta_{ab}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&ra(t)\left(r\dot{a}+1\right)&0\\ 0&0&0&ra(t)\sin^{2}\theta\left(r\dot{a}+1\right)\end{array}\right),$ (C.46)
leading to the expansion scalar
$\theta=q^{ab}\theta_{ab}=\frac{2(r\dot{a}+1)}{ra(t)}.$ (C.47)
The shear tensor vanishes identically,
$\sigma_{ab}=0,$ (C.48)
reflecting the homogeneity and isotropy of the FLRW spacetime.
The null Brown–York tensor on the cosmological null boundary is obtained as
$T^{a}{}_{b}\big|_{\rm(null)}=\left(\begin{array}[]{cccc}-\frac{r\dot{a}+1}{8\pi ra(t)}&-\frac{r\dot{a}+1}{8\pi r}&0&0\\ -\frac{r\dot{a}+1}{8\pi ra(t)^{2}}&-\frac{r\dot{a}+1}{8\pi ra(t)}&0&0\\ 0&0&-\frac{2r\dot{a}+1}{8\pi ra(t)}&0\\ 0&0&0&-\frac{2r\dot{a}+1}{8\pi ra(t)}\end{array}\right).$ (C.53)
The energy density measured by the null observer is therefore
$\epsilon_{\rm null}=-\frac{1+r\dot{a}}{4\pi ra(t)},$ (C.54)
while the momentum density vanishes, $p^{c}=0$. The spatial stress tensor is given by
$s^{ab}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&-\frac{2r\dot{a}+1}{8\pi r^{3}a^{3}(t)}&0\\ 0&0&0&-\frac{(2r\dot{a}+1)\csc^{2}\theta}{8\pi r^{3}a^{3}(t)}\end{array}\right).$ (C.59)
Appendix D Energy conservation relation in null hydrodynamics